### v1

# Searching for MeV-scale Axion-like Particles and Dark Photons with PandaX-4T Tao Li

Searching for MeV-scale Axion-like Particles and Dark Photons with PandaX-4T
Tao Li
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Zihao Bo
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Wei Chen
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xun Chen
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yunhua Chen
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Zhaokan Cheng
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Xiangyi Cui
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Yingjie Fan
Department of Physics, Yantai University, Yantai 264005, China
Deqing Fang
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Zhixing Gao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Lisheng Geng
School of Physics, Beihang University, Beijing 102206, China
Peng Huanwu Collaborative Center for Research and Education, Beihang University, Beijing 100191, China
International Research Center for Nuclei and Particles in the Cosmos & Beijing Key Laboratory of Advanced Nuclear Materials and Physics, Beihang University, Beijing 100191, China
Southern Center for Nuclear-Science Theory (SCNT), Institute of Modern Physics, Chinese Academy of Sciences, Huizhou 516000, China
Karl Giboni
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Xunan Guo
School of Physics, Beihang University, Beijing 102206, China
Xuyuan Guo
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Zichao Guo
School of Physics, Beihang University, Beijing 102206, China
Chencheng Han
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Ke Han
ke.han@sjtu.edu.cnSchool of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Changda He
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinrong He
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Di Huang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Houqi Huang
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Junting Huang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Ruquan Hou
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yu Hou
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xiangdong Ji
Department of Physics, University of Maryland, College Park, Maryland 20742, USA
Xiangpan Ji
School of Physics, Nankai University, Tianjin 300071, China
Yonglin Ju
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Chenxiang Li
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jiafu Li
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Mingchuan Li
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Shuaijie Li
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Zhiyuan Li
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Qing Lin
State Key Laboratory of Particle Detection and Electronics, University of Science and Technology of China, Hefei 230026, China
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Jianglai Liu
jianglai.liu@sjtu.edu.cnNew Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Congcong Lu
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xiaoying Lu
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Lingyin Luo
School of Physics, Peking University, Beijing 100871, China
Yunyang Luo
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Wenbo Ma
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Yugang Ma
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Yajun Mao
School of Physics, Peking University, Beijing 100871, China
Yue Meng
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Xuyang Ning
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Binyu Pang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Ningchun Qi
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Zhicheng Qian
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiangxiang Ren
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Dong Shan
School of Physics, Nankai University, Tianjin 300071, China
Xiaofeng Shang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiyuan Shao
School of Physics, Nankai University, Tianjin 300071, China
Guofang Shen
School of Physics, Beihang University, Beijing 102206, China
Manbin Shen
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Wenliang Sun
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yi Tao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Anqing Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Guanbo Wang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Hao Wang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jiamin Wang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Lei Wang
College of Nuclear Technology and Automation Engineering, Chengdu University of Technology, Chengdu 610059, China
Meng Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Qiuhong Wang
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Shaobo Wang
shaobo.wang@sjtu.edu.cnSchool of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Siguang Wang
School of Physics, Peking University, Beijing 100871, China
Wei Wang
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Xiuli Wang
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xu Wang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Zhou Wang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yuehuan Wei
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Weihao Wu
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yuan Wu
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Mengjiao Xiao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiang Xiao
xiaox93@mail.sysu.edu.cnSchool of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Kaizhi Xiong
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yifan Xu
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Shunyu Yao
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Binbin Yan
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Xiyu Yan
School of Physics and Astronomy, Sun Yat-Sen University, Zhuhai 519082, China
Yong Yang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Peihua Ye
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Chunxu Yu
School of Physics, Nankai University, Tianjin 300071, China
Ying Yuan
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Zhe Yuan
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Youhui Yun
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xinning Zeng
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Minzhen Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Peng Zhang
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Shibo Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Shu Zhang
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Tao Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Wei Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Yang Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Yingxin Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Yuanyuan Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Li Zhao
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Jifang Zhou
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Jiaxu Zhou
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Jiayi Zhou
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Ning Zhou
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Xiaopeng Zhou
School of Physics, Beihang University, Beijing 102206, China
Yubo Zhou
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Zhizhen Zhou
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
(September 1, 2024)
Abstract
Axion-like particles (ALPs) and dark photons (DPs) are viable dark matter particle candidates.
We have searched for possible ALP/DP signals in the PandaX-4T liquid xenon detector using 94.8 days of data.
A binned likelihood fit is constructed to search for possible mono-energetic peaks induced by the absorption processes between ALPs/DPs and atomic electrons of xenon.
A detailed temporal model of decays associated with xenon isotopes is introduced to constrain the number of background events.
No signal excess over background expectations is observed, and we have established the most stringent exclusion limits for most ALP/DP masses ranging from 150 keV/ to 1 MeV/.
Astronomical and cosmological observations have provided compelling evidence for the existence of dark matter (DM) [1, 2, 3], which is crucial for understanding the evolution of the universe.
For the last few decades, many terrestrial experiments worldwide have been dedicated to the search for DM, with a particular emphasis on Weakly Interacting Massive Particles (WIMPs) [4, 5, 6, 7], a prevailing candidate of cold dark matter (CDM).
However, no conclusive signals from WIMPs have been detected so far.
On the other hand, there are some anomalies observed in the small-scale structure in galaxies which seem inconsistent with simulations within the CDM framework [8, 9, 10, 11, 12].
This has prompted increased interest in alternative models involving lighter DM particles with weaker couplings to Standard Model (SM) particles [13, 14, 15].
Among these models, the axion-like particles (ALPs) and dark photons (DPs), also referred to as bosonic super-WIMPs [16, 17, 18], are of experimental interest.
In contrast with the elastic scattering of WIMP with nucleus or electrons, they can be searched via unique absorption signals [18].
For example, the axioelectric effect, analogous to the photoelectric effect, will lead to the absorption of ALPs by the detector target with the energy transferred to one of the atomic electrons, producing a mono-energetic signal at the rest mass of ALPs.
The absorption cross-section equals to [18],
in which () is the rest mass of the ALP (electron), is the fine-structure constant, is the velocity of the incoming ALP, is the dimensionless coupling constant between the electrons and ALP, and is the photoelectric effect cross-section for a photon with an energy of .
Similarly, the cross-section for DP is
where is the kinetic mixing constant between the DP and the real photon.
Assuming that ALPs or DPs consist of all the DM in our galaxy with a density of 0.3 GeV/cm3, the corresponding event rate in a terrestrial detector can be obtained as
respectively, where represents the atomic mass of the absorbing atoms in the detector.
The ALP (DP) mass () is in the unit of keV/ and in the unit of barn.
The constants and are measured in experiments.
Among the searches of ALPs and DPs [19, 20, 21, 22, 23, 24, 25, 26, 27, 28], XENONnT [20] is leading the limit at the masses below 140 keV/, while GERDA [21] and COSINE-100 [22] have set the most significant constraints in the
(100) keV/ to 1 MeV/ range.
In this paper, we use the commissioning data set (Run0) of the PandaX-4T experiment, a total of 94.8 days of data from November 28, 2020, to April 16, 2021, to search for ALP and DP signals.
The targeted ALP or DP masses are between 30 keV/ and 1 MeV/, while the search is performed in an energy region of interest (ROI) of 25 keV to 1 MeV.
Compared to previous analyses of PandaX-4T in the MeV energy range [29, 30], the energy reconstruction procedure is further optimized to improve the energy resolution.
The time-varying background contributions from short-lived xenon isotopes, including 127Xe, Xe, and Xe, are now incorporated into the modeling for the first time in PandaX-4T.
Furthermore, we have developed a convolution method to propagate uncertainties of energy response into the energy spectrum to fully incorporate detector uncertainties in the likelihood fit.
The PandaX-4T detector is a cylindrical, dual-phase time projection chamber (TPC) measuring 118.5 cm in diameter and 118.5 cm in height.
The active volume containing 3.7 tonnes of natural xenon is surrounded by a field cage with an anode on the top and a cathode on the bottom.
Two three-inch Hamamatsu PMT arrays are installed above the anode and below the cathode for signal readout.
A detailed description of the detector can be found in Ref. [31].
The detector measures the energy deposition and its three-dimensional position via the scintillation signal () and the electroluminescence signal (), which scales with number of ionized electrons.
The data production and event selection procedures are similar to Refs. [29, 30, 32, 33], where the reconstruction of single-site (SS) spectrum from 25 keV to 2.8 MeV is achieved.
The total energy of an event is calculated by combining and collected by the bottom PMT array, according to the formula [34].
PDE, EEE, and are the photon detection efficiency for , electron extraction efficiency, and the single-electron gain for , respectively.
The horizontal position is obtained based on the maximum likelihood estimation with the observed charge distribution in the top PMT array and the photon acceptance functions derived from optical Monte Carlo (MC) simulations.
We have optimized the position reconstruction in the vertical (), radial () and azimuthal () directions using calibration data from Kr and wall events from 210Po particles.
This determines the fiducial mass (FM) to be kg by scaling the percentage of [30], corresponding to the total exposure of kg yr.
The detector is calibrated using external radioactive 137Cs, 60Co, and 232Th sources.
The non-uniformity of spatial energy response is monitored and corrected using the internal calibration source Kr.
Compared with the previous analysis [30], the energy reconstruction has been improved in three aspects.
The temporal variation of energy response is characterized by the signals of 222Rn progenies and corrected accordingly.
The linearity of energy reconstruction is further improved by adding one more high-statistics 41.5 keV peak from Kr, in additional to Xe, Xe, 127Xe, 60Co, 40K, and 232Th.
Consequently, we have improved the energy resolution, specifically from 3.6% to 3.0% at 208 keV, which allows improvements in estimating activities of short-lived xenon isotopes mentioned later.
Table 1:
Summary of sources of systematic uncertainties.
represents the 5-parameter detector response model (see text), with means and uncertainties determined from calibration fit.
The energy response is modeled with five parameters.
The energy resolution is modeled as a Gaussian function with the width constructed as , with energy in the unit of keV.
The energy scale is defined as to account for possible bias with respect to the reconstructed energy .
The measured energy spectrum is a convolution of the true energy spectrum with the five-parameter response model.
The parameters and uncertainties are determined by fitting the peaks of 41.5 keV (Kr), 164 keV (Xe), 236 keV (Xe and Xe) from the calibration data, and 1460 keV (40K) outside the ROI from Run0, therefore completely uncorrelated with the final search fit.
The extracted values and uncertainties of the parameters (Table 1) will be used as priors, together with the covariance matrix in fitting the Run0 data.
The total detection efficiencies for signals of ALPs/DPs are the product of data quality cut efficiency, SS cut efficiency, and ROI acceptance.
Quality cut variables, used to eliminate noise and select electronic recoil events, are adopted from Ref. [30] but have been adjusted to account for events down to 25 keV.
The adjusted cut criteria are validated using calibration data and subsequently applied to the entire Run0, resulting in an efficiency of ()%.
The identification of SS and multi-site (MS) events follows the method in Ref. [29].
Charge deposits in a given event may be separated into different clusters, allowing classification as SS or MS based on the number of observed peaks.
The ratio of the number of SS and SS+MS events within the ROI is calculated using BambooMC, a Geant4-based Monte Carlo (MC) framework [35], and verified through 232Th calibration data.
The SS/(SS+MS) ratio of 232Th calibration data in the ROI is %, with the uncertainty from the difference between data and simulation averaged over the energy range.
The relative systematic uncertainty from 232Th calibration data is conservatively used for both the signal and background for the SS fraction.
On top of the SS fraction, the quality cut efficiency is energy dependent, ranging from 100.0% to 95.7% in the ROI based on MC.
The ROI acceptance for signals is close to 100%.
The liquid xenon density and FM are two other systematic factors for ALP/DP detections.
The liquid xenon density is inherited from Ref. [29] with a relative uncertainty of 0.13%.
The relative difference between the FM calculated from geometry and that derived from Kr events distribution is 1.57%, which is taken as the systematic uncertainty.
In our ROI, the background contribution originates from the detector materials, liquid xenon, and solar neutrinos, as shown in Table 2.
The activities of 232Th, 238U, 60Co, and 40K in detector materials have been reported in Ref. [29].
The expected events in ROI are , , , and , respectively.
The concentration of Kr is determined by - cascades through the metastable state Rb.
The Kr/Xe concentration is parts per trillion, assuming an isotopic abundance of for Kr [36].
The 212Pb activity of is determined from the rate of 212Po [30].
The 214Pb rate is left float in the fit, the same as the previous analyses [29, 30].
The energy spectrum of the elastic scattering of solar and 7Be neutrinos on electrons is adopted from the Ref. [37], with an uncertainty of approximately 10% [38], resulting in events within ROI.
Figure 1: Time evolution of 127Xe, Xe and Xe in SS spectrum during the Run0 of PandaX-4T experiment.
The starting time of the axis is defined as the beginning of the dataset.
AmBe and PuC refer to sources used in the neutron calibrations.
Table 2: The background contribution in the ROI. The fitted counts are from the background only fit.
Components
Expected
Fitted
232Th
238U
60Co
40K
85Kr
212Pb
214Pb
136Xe
124Xe
125Xe
125I
133Xe
164 keV
208 keV
236 keV
380 keV
408 keV
Be
Figure 2: Fit of background and a hypothetical 990 keV/ DM signal to SS data from 25 keV to 1050 keV with bin size of 1 keV.
The x-axis represents the reconstructed energy in the data.
Xe∗ includes the contributions from 124Xe, 125Xe, 127Xe, Xe, Xe, and 133Xe.
The lower panel shows the residuals together with () bands.
The contributions from 136Xe and 124Xe double electron capture are calculated based on the half-life measurements in PandaX-4T [29, 39].
The expected events in ROI are and , respectively.
Short-lived xenon isotopes, Xe, Xe, Xe and Xe were induced by neutron calibration.
Xe and Xe results in mono-energetic peaks at 164 keV and 236 keV, with the half-life of 11.8 days and 8.9 days, respectively.
125Xe, with a half-life of 16.9 hours, undergoes electron capture and decays to the relatively long-lived isotope 125I, with a half-life of 59.4 days.
Given the low statistics, 125Xe is treated as a floating parameter in the fit.
The activity of 125I is calculated based on its temporal modeling [39], resulting in events in the ROI.
133Xe is also left float in the fit since the spectrum is continuous, characterized by a 346 keV endpoint decay with corresponding de-excitation 81 keV ray.
In addition, 127Xe was introduced by a batch of approximately 30 kg of xenon from above-ground (exposed to cosmogenic neutrons), injected into the detector during the data taking.
127Xe decays via electron capture with a half-life of 36.4 days.
The de-excitation of the daughter 127I generates mono-energetic peaks at 208 keV, 236 keV, 380 keV, and 408 keV.
The estimated numbers of 127Xe, Xe, and Xe events are obtained by fitting the time evolution in the SS data, with decay half-lives also set free.
In order to increase statistics, a larger FM of 2.43 tonnes is selected.
Each peak is characterized by a Gaussian function plus a linear background.
The evolution of Xe is fitted as a single component, while the evolution of four peaks of Xe and Xe is fitted simultaneously to decompose the contribution of two isotopes at 236 keV.
The resulting time evolution and the corresponding measured half-lives are shown in Fig. 1, with fitted half-lives consistent with existing values in nuclear databases [40].
Based on the temporal fit, the estimated number of events for the 164 keV, 208 keV, 236 keV, 380 keV, and 408 keV peaks in ROI are shown in Table 2.
We employ a binned likelihood method, with the likelihood function constructed as
where and are the expected and observed numbers of events in the energy bin, respectively.
The Gaussian penalty term of the energy response contains the five-parameter (Table 1) and the covariant matrix .
The Gaussian penalty terms constrain the nuisance parameters , , and , which are the relative uncertainties of the overall efficiency, the signal selection (Table 1), and the background model (Table 2), respectively.
is the number of Gaussian-constrained nuisance paramters.
is defined as
where and are the counts of signal and background component , respectively.
The corresponding and are the bin values of the normalized energy spectrum convolved with the five-parameter energy response model.
A background-only fit is performed before the signal fits, with /NDF=1.14.
The fitted background contributions are shown in Table 2 and agree with the expected values.
Background is dominated by short-lived xenon isotopes, 136Xe 2, 214Pb decay, and detector material.
Fitted nuisance parameters are within 1 of the input values, except for and in the energy response model, which are pulled by 1.9 and 2.5 , respectively.
This suggests that the energy resolution function derived solely by mono-energetic peaks is insufficient to describe the full measured spectrum.
We conduct a scanned fit to the SS spectrum, including Gaussian peaks of the hypothetical DM signals, with DM masses spanning from 30 keV/ to 1 MeV/ with a step size of 10 keV/.
An illustrative example of the fit results for DM mass of 990 keV/ with data-fit residuals is presented in Fig. 2.
At this mass, the majority of residuals is less than 3.
The fitted signal count is .
The local significances of two DM masses, 250 keV and 260 keV, are found to be between 3 and 4, occurring near the Gaussian background of 236 keV (Xe and Xe).
Taking into account the look elsewhere effect [41, 42], the global significance is less than 1. Therefore, no significant evidence for a signal is observed within the mass range of [30 keV/, 1 MeV/].
Figure 3:
The 90% C.L. upper limits of the couplings of ALPs (top) and DPs (bottom) to atomic electrons.
The results from this work are represented by the red curve starting at 30 keV, while the curve below 25 keV is taken from Ref. [43].
The upper limits at 90% confidence level (C.L.) on the event rate have been set and converted to upper limits of coupling strength (Fig. 3), as described in Eq. 1.
Leading direct detection limits from other experiments are also plotted for comparison [26, 20, 22, 21].
The relative deterioration at certain masses in the limit curves is due to background fluctuations, such as the mono-energetic background at 164 keV from Xe.
Our limits are the most competitive over a wide region of the masses, ranging from 150 keV/ to 1 MeV/, with an average improvement of 1.5 times better than the existing results [22, 21].
The coupling of ALPs (DPs) to electrons of () at the mass of 990 keV/ derived from the fit shown in Fig. 2 is 2.5 times better than the COSINE-100 result [22].
Compared to other experiments, the improvement is due to a combination of large exposure, low background rate, and broader energy range.
It is noteworthy that the weakening of our upper limits with increasing mass is primarily due to the steep decrease of the photoelectric cross-section, as we solely searched for absorption events.
In the MeV mass regions of the ALPs and DPs, the cross-section of Compton-like process [22], which are mostly MS events, becomes more prominent.
A dedicated analysis focused on MS events is underway to improve the signal detection efficiency.
In summary, we have searched for ALPs and DPs with masses up to 1 MeV/ using 162.3 kg yr exposure of PandaX-4T Run0.
A detailed analysis of the time evolution of xenon isotopes improves the background modeling, and the inclusion of energy response model convolution in the likelihood function results in a more rigorous treatment of systematic uncertainties.
No significant excess over the expected background is observed, and the upper limits at 90% C.L. on the effective couplings between ALPs/DPs and atomic electrons of xenon are derived.
Our limit is comparable with other direct searches in the DM mass range from 30 to 150 keV/ and the most competitive over a large mass range from 150 keV/ to 1 MeV/.
Between 2021-2022, PandaX-4T completed the first science run (Run 1) with 163.5 days data taking, and has recently completed an upgrade and resumed physics data taking.
Larger statistics and higher-quality data can further enhance the understanding of the background model and the search sensitivities of the ALPs and DPs.
This project is supported in part by grants from National Science Foundation of China (Nos. 12090060, 12090061, 12305121, U23B2070), a grant from the Ministry of Science and Technology of China (Nos. 2023YFA1606200, 2023YFA1606201, 2023YFA1606202), China Postdoctoral Science Foundation (No. 2023M744093), and Office of Science and Technology, Shanghai Municipal Government (grant No. 22JC1410100, 21TQ1400218).
We thank for the support by the Fundamental Research Funds for the Central Universities. We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Hongwen Foundation in Hong Kong, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.